# Robust Weight Imprinting: Insights from Neural Collapse and Proxy-Based Aggregation

### v2

@@ Abstract
The capacity of a foundation model allows for adaptation to new downstream tasks. Weight imprinting is a universal and efficient method to fulfill this purpose. It has been reinvented several times, but it has not been systematically studied. In this paper, we propose a framework for imprinting, identifying three main components: generation, normalization, and aggregation. This allows us to conduct an in-depth analysis of imprinting and a comparison of the existing work. We reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization. We determine proxies through clustering and propose a novel variant of imprinting that outperforms previous work. We motivate this by the neural collapse phenomenon – an important connection that we can draw for the first time. Our results show an increase of up to 4% in challenging scenarios with complex data distributions for new classes. Finally, we publicly release our code at https://github.com/DATEXIS/multi-imprinting/.
@@ 1 Introduction
In machine learning applications, training models from scratch is often not viable due to limitations in data and compute. A popular solution is to apply transfer learning [1, 2] based on foundation models (FMs) [3] that are pre-trained on a large amount of data. A common approach in practice to adapt an FM to a novel task is to freeze its parameters and replace the output layer with a new head, e.g., for classification.
@@
Qi et al. [4] propose a simple solution for few-shot classification, called imprinting. Namely, the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean. These class means are representatives of the classes, which we generally call proxies. This results in an efficient method without the need for gradient-based optimization. A plethora of studies have emerged surveying this technique by adding complexity and adaptability [5, 6, 7, 8, 9, 10, 11]. Despite many adaptations, imprinting lacks a systematic comparison that unifies them. Understanding its variations could unlock greater efficiency and performance across many fields, making the method even more versatile and impactful.
@@
We present a unifying framework that enables a systematic comparison of existing imprinting techniques. More precisely, we generalize prior work by decomposing imprinting into three principal steps (see fig. 1). During generation (GEN) of weights, the network selects representative data samples and generates one or more weight vectors per class (proxies). Normalization (NORM) is crucial, as the network needs to balance its generated weight vectors. Aggregation (AGG) entails the computation of the final output, e.g., a class label.
The efficiency of imprinting enables us to perform a comprehensive analysis with a large number of experiments. We present a new, best-performing imprinting strategy using multi-modal weight imprinting in combination with the correct way of normalization, outperforming previously studied methods, as depicted in fig. 2.
@@
We investigate a recently discovered phenomenon called neural collapse [12], which provides a compelling explanation for why imprinting works. According to this phenomenon, when neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means [12, 13]. Our investigation proves that a measurement of neural collapse provides insights about imprinting.
@@
  • We deconstruct weight imprinting into a framework composed of generation, normalization, and aggregation, and discuss variations for each of them, identifying prior work as special cases (section 3). To the best of our knowledge, we are the first to conduct a comprehensive analysis of imprinting to this scale (section 5).
  • We present a new imprinting method utilizing $k$-means clustering for weight generation (section 6.1) and show its benefits in certain few-shot scenarios (section 6.2).
  • To the best of our knowledge, we are the first to identify a connection between imprinting success and measures of neural collapse (section 6.3).
We make the source code to reproduce our results publicly available. [^1] https://github.com/DATEXIS/multi-imprinting/
Figure 1: Overview of our imprinting framework. The foundation model FM is frozen and shows neural collapse. The weight generator (GEN) uses training data from a novel task T to generate one or more weight vectors (proxies) per class $1,\dotsc,c$. The final output for the test data in T is computed by an aggregation (AGG) mechanism. Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively. During inference, embeddings are normalized according to NORMinf (not shown here).
Paper NORMpre GEN NORMpost NORMinf AGG Avg. acc. %
Qi et al. [4] L2 mean L2 L2 max 87.72
Hosoda et al. [14] none mean quantile none max 80.71
Janson et al. [15] none mean none none 1-nn 87.64
Ours L2 k-means L2 L2 max 91.48
@@
[table]A table beside a figure
Figure 2: Previously studied imprinting strategies are special cases within our framework. The framework enables the creation of a novel configuration ("Ours") that outperforms previous work across FMs and Ts by a large margin with statistical significance. Here, $k=20$ is chosen.
@@ 2 Related Work
Imprinting and Few-Shot Learning.
Weight imprinting was introduced in [4] for the few-shot learning scenario. It is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples. Qi et al. [4] find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms. However, we are not limiting the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency. Imprinting has also been applied to object detection [5, 6], multi-label classification [8], semantic segmentation [10], and in combination with an attention mechanism to generate weights for the novel classes in a few-shot classification task [11].
Hosoda et al. [14] apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights. We consider this as one normalization scheme in our framework. Zhang et al. [16] apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available. They speculate whether normalization is a constraint in their imprinting model.
Before the era of deep learning, Mensink et al. [17] analyze the transferability of hand-crafted image features. They use a "nearest class multiple centroids" (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm. In combination with metric learning, they compare favorably against the $m$-nearest neighbor algorithm. Our work, on the other hand, highlights efficient transfer learning provided by foundation models.
@@
Our work is related to the use of embedding vectors extracted with pre-trained models, which is one of the most straightforward transfer learning techniques, since the seminal works in computer vision [18] and natural language processing [19]. Kornblith et al. [20] showed that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks. In addition, Huh et al. [21] provided insights into the required quality of pre-training data. Our work is orthogonal to these studies, since we focus on studying weight generation, normalization, and aggregation techniques applied later on for new task adaptation.
@@
Class means have also been used as proxies in CL. Although we investigate transfer learning scenarios, we review the imprinting applications and results from CL. Rebuffi et al. [22] dynamically select a subset of examples for each class and update internal representations via gradient descent. They use a nearest mean classifier (NMC) with respect to the saved examples. Janson et al. [15] use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings. However, they do not investigate the effect of normalization and using multiple proxies.
Findings of [23] show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored. In our work, however, we compare imprinting all data to a limited number of more representative proxies striving for efficiency.
@@
The phenomenon of NC was identified by [12] and refers to the convergence of the last-layer weight vectors to class means. It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) [13, 24, 25], but complete neural collapse is practically unrealistic [26]. In transfer learning, Galanti et al. [27] show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios. In our work, we expand the survey on NC by experimenting with out-of-distribution classes belonging to different datasets and linking their degree of collapse to the success of certain imprinting strategies.
3 Imprinting Framework
In order to find out how to best set the classifier weights of a foundational model in downstream tasks T, we create a framework (see fig. 1) that encompasses many different combinations, all of which work without gradient-based training. Thereby, we can unify all the existing imprinting strategies described in section 2.
We analyze multi-class classification scenarios in that we do not separate into base and new classes, but focus on all classes in a novel T at the same time. To investigate the effect of the number of samples given, we look at n-shot ($n\in\mathbb{N}$) scenarios. For that, we randomly pre-sample the training data of T to $n$ samples per class – transitioning into the regime of few-shot learning.
@@ 3 Imprinting Framework
We analyze the effect of weight generation (GEN), normalizations ($\texttt{NORM}=\{\texttt{NORM}\textsubscript{{pre}},\texttt{NORM}\textsubscript{{post}},\texttt{NORM}\textsubscript{{inf}}\}$), and aggregation (AGG). The framework depicted in fig. 1 consists of three main building blocks: a foundation model FM, a weight generator GEN, and extendable classifier weights that are imprinted. The FM remains frozen throughout the experiments. It receives data from T as inputs and produces embedding vectors. The training process generates weight vectors for each of the $c$ classes in T. Hereby, embeddings from the FM are normalized before the generation (GEN) step according to NORMpre. The generated weight vectors per class are called proxies, prototypes, or representatives [28, 29, 30, 31]. These proxies are normalized according to NORMpost. As in [4], we do not use bias values. To classify the test data in T during inference, it is first embedded by the FM, normalized according to NORMinf, and finally aggregated by AGG, resulting in a predicted class label.
Special Imprinting Cases.
Previously proposed imprinting methods can be defined as a special case of our framework. Figure 2 in section 6.1 provides an overview of existing imprinting strategies listed in literature for foundational models and benchmarks them with the new best-performing one we find through our framework. In total, we inspect all possible combinations (including variations in models, tasks, and seeds).
@@
The purpose of GEN is to determine how the embeddings of the training data in T are used to form the new weights. In contrast to [4] which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in [17]. This enables non-linear classification. We denote the number of proxies as $k$, ranging between $1$ and the number of samples. We investigate the following operations conducted per class to generate its proxies:
@@
  • k-means: $k$-means cluster centers. $k=1$ is the same as mean.
  • k-medoids: $k$-medoids cluster centers.
  • k-cov-max (covariance-maximization): Top $k$ embeddings by covariance.
  • k-fps (farthest-point sampling): Iteratively selecting $k$ embeddings, such that it maximizes the distance from already selected ones (starting with random sample).
We choose this diverse list of methods to cover a wide range of approaches, ranging from heuristics (e.g., k-fps) to more complex algorithms (e.g., k-means). Note that only mean and k-means generate proxies that are not included in the given samples.
@@
L2 normalization can be applied to embeddings before GEN via NORMpre, to the generated weights via NORMpost, and to embeddings in inference via NORMinf. In any case, the vector is $L^{2}$-normalized by dividing it by its Euclidean $\|\cdot\|_{2}$ length.
quantile normalization [32, 33] can only be applied to generated weights. This non-linear operation distributes weights equally. Recall that if more than one class is contained in T ($c>1$), GEN is performed for each class, and the reference distribution changes accordingly. In particular, for the first class there is no reference distribution to map to. This is different from [14], where new weights are matched to the distribution of the original classifier weights of the FM. Since we do not consider the classes used for pre-training the FM and especially do not assume access to their last-layer weights, this is not possible in our scenario.
@@
There are various ways to use the generated weights per class during inference, especially when $k>1$. We focus on two different modes, max and m-nn. The former, max, computes the inner product of the input embedding and the imprinted weights and outputs the class label with the maximum activation. The latter, m-nn, uses the class weights as keys and the embeddings as values, and chooses the final winning output class via the $m$-nearest neighbor algorithm. The m-nn voting is weighted by the inverse of the distances to their nearest neighbor, turning it into weighted majority voting.
Note that max is the same as 1-nn in the case of L2 for NORMpost, since for any fixed embedding vector $v$ and variable proxy $w$, the argmin of $\|v-w\|^{2}=\|v\|^{2}-2\langle v,w\rangle+\|w\|^{2}$, calculated by 1-nn, is the same as the argmax of the inner product $\langle v,w\rangle$ calculated in max.
4 Measurement of Neural Collapse
Neural collapse (NC) [12] refers to the phenomenon that occurs on the last-layer classifier weights of neural networks in the terminal phase of training (TPT). When the network is trained well beyond zero training error, the learned embeddings of each class, assuming balanced classes, collapse to their class means. These globally centered class means and classifier weights form a simplex equiangular tight frame (ETF) – a collection of equal length and maximally equiangular vectors, that maximize the between-class variability. This results in an optimal linearly separable state for classification. In fig. 3 (left), we illustrate the collapse of a FM on its pre-training data. The newly arrived data T from a different dataset is distributed more unevenly across the embedding space (right).
Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification. We focus on variability collapse ($\mathcal{NC}_{1}$) as in [13]:
$\mathcal{NC}_{1}=\tfrac{1}{c}\mathrm{trace}(\Sigma_{W}{\Sigma^{+}_{B}}),$ (1)
where $\Sigma_{W},\Sigma_{B}\in\mathbb{R}^{l\times l}$ are within- and between-class covariance matrices, respectively, $l$ is the dimension of the embedding vector, $c$ is the number of classes, and + symbolizes the pseudo-inverse. Based on the equation, an $\mathcal{NC}_{1}$ score closer to zero signifies a higher collapse. In contrast, an increase in multi-modality of data leads to a higher $\mathcal{NC}_{1}$ score (as analyzed in fig. 11). Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $c$. According to NC, imprinting the mean, as originally done in [4], is best when $\mathcal{NC}_{1}$ is small. We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class. We investigate this in section 6.3.

### v3

@@ Abstract
The capacity of foundation models allows for their application to new, unseen tasks. The adaptation to such tasks is called transfer learning. An efficient transfer learning method that circumvents parameter optimization is imprinting. The conceptual differences between studies on imprinting form the basis of our systematic investigation. In this work, we propose the general IMPRINT framework, identifying three main components: generation, normalization, and aggregation. Through the lens of this framework, we conduct an in-depth analysis and a comparison of the existing methods. Our findings reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization. Beyond an extensive analytical grounding, our framework enables us to propose a novel variant of imprinting which outperforms previous work on transfer learning tasks by 4%. This variant determines proxies through clustering motivated by the neural collapse phenomenon – a connection that we draw for the first time. We publicly release our code at https://github.com/DATEXIS/IMPRINT.
@@ 1 Introduction
In machine learning applications, training models from scratch is often not viable due to limitations in data and compute. A popular solution is to apply transfer learning (transferlearning3; transferlearning2) based on foundation models (FMs) (foundationmodels) that are pre-trained on a large amount of data. To tune a FM to a novel task, e.g., for classification, a common procedure is to freeze the model parameters and replace the output layer with a new head. A particularly simple method for implementing such a new head was proposed by qi-imprinting and coined imprinting.
@@
In the original work by qi-imprinting, the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean. These class means are representatives of the classes, which we denote as proxies. In general, we refer to imprinting as efficient learning methods without the need for cross-class statistics or gradient-based optimization. A plethora of studies have emerged surveying this technique by adding complexity and adaptability (fewshot-metagenweights; masked-imprinting; hypersphere-imprinting; imprint-multilabel; imprint-objdetec; multiple-projection-head-imprinting; yan2023few). Despite its many adaptations, imprinting lacks a systematic comparison that unifies them. Understanding its variations could unlock greater efficiency across many fields, making the method even more versatile and impactful.
Applications.
In particular, while in some practical applications, accuracy is prioritized over computational efficiency, the latter does become a critical requirement in scenarios where computational resources are severely constrained, e.g., in the chemical and polymer processing industries. Here, battery-powered edge compute is essential and frequent retraining or large-scale optimization is infeasible. Imprinting has proven particularly effective in these contexts of edge-embedded devices (hypersphere-imprinting). For instance, imprintingrobot implement a vision-based robotic force grasping with a variable-stiffness gripper that can safely handle both fragile and heavy objects by rapidly adapting to novel categories without retraining (Continual Learning (CL) setting). Industrial adoption also underscores this trend with Google’s Coral Edge TPU including an ImprintingEngine API (coralai), that allows users to add new classes from a few examples without recompiling the model. Very recently, (belal2025fsid) apply imprinting to spectrogram embeddings from IMU/EMG gait data, achieving efficient classification in low-data Human Activity Recognition (HAR) tasks. Apart from that, ncmbaseline extend imprinting to CL, establishing a simple baseline with competitive performance compared to more sophisticated state-of-the-art CL algorithms.
Figure 1: Overview of our IMPRINT framework. The foundation model FM is frozen and shows neural collapse. The weight generator (GEN) uses training data from a novel task T to consecutively generate one or more weight vectors (proxies) per class $1,\dotsc,C$ in T. In inference, the final output for the test data in T is computed by an aggregation (AGG) mechanism. Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively. During inference, embeddings are normalized according to NORMinf.
@@
We present IMPRINT, a framework that enables a comprehensive analysis of existing imprinting techniques. More precisely, we generalize prior work by decomposing imprinting into three principal steps (see fig.˜1). During generation (GEN) of weights, the method selects representative data samples and generates one or more weight vectors (proxies) per class. Normalization (NORM) is crucial, as the generated weights need to be balanced. Aggregation (AGG) entails the computation of the final output, e.g., a class label. The computational efficiency of imprinting allows us to perform a large number of experiments. Through IMPRINT, we are able to propose a novel, best-performing imprinting strategy using multi-proxy weight imprinting in combination with $L^{2}$ normalization, outperforming previously studied methods, as depicted in fig.˜2.
@@
When neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means (papyan2020prevalence; zhu2021geometric). We investigate this phenomenon as a potential explanation for when and why imprinting works. Our analysis proves that there exists a relationship between a measure of neural collapse and the success of imprinting. Since quantification of this phenomenon is possible in a post-hoc fashion over the FM features, we believe that these insights will contribute to further development of imprinting methods, as well as, training regimes of FMs that are more suitable for transfer learning via imprinting.
@@
  • We deconstruct weight imprinting into the IMPRINT framework composed of generation, normalization, and aggregation, and discuss variations for each of them, while identifying prior work as special cases (section˜3.1). To the best of our knowledge, we are the first to conduct a comprehensive analysis of imprinting to this scale (section˜4).
  • We present a new, outperforming imprinting method utilizing $k$-means clustering for weight generation (section˜5.1) and show its benefits in certain low-data regimes (section˜5.2).
  • As far as we are aware, we are the first to identify a relationship between the degree of neural collapse and imprinting success (section˜5.3).
@@
[…]
Figure 2: Previously studied imprinting strategies are special cases within IMPRINT. We evaluate 12 different classification tasks Ts derived from MNIST, FashionMNIST, and CIFAR-10, each with 10 classes or subsets thereof, and 4 pre-trained models FMs (resnet18, resnet50, vit_b_16, swin_b). The proposed configuration (“Ours”) derived from IMPRINT outperforms previous work across FMs and Ts by a large margin with statistical significance, as confirmed by the critical difference (CD) diagram below. Since absolute accuracies vary substantially across tasks and models (as reflected by the large standard deviations (std)), this rank-based aggregation is used for fair comparison. Here, $k=20$ is used, highlighting the gain of using multiple proxies per class. For reference, the gray row reports an oracle method that uses cross-class feature statistics to generate weights (see section˜A.5). It is not an imprinting method and therefore not directly comparable to the imprinting-based approaches above. Nonetheless, the results indicate that our method substantially narrows the gap between single-proxy mean imprinting and this oracle baseline.
@@ 2 Related Work
Imprinting and Low-Data Regimes.
Weight imprinting is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples and was first introduced by qi-imprinting for the few-shot learning scenario. The authors find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms. In contract, in our work we do not limit the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency.
Imprinting has been applied to object detection (imprint-objdetec; yan2023few), multi-label classification (imprint-multilabel), semantic segmentation (masked-imprinting), and combined with an attention mechanism to generate weights for the novel classes in a few-shot classification task (fewshot-metagenweights).
itsdone apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights. We consider this as one normalization scheme in our framework. covidimprint apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available. They speculate whether normalization is a constraint in their imprinting model.
Before the era of deep learning, first-nearestclassmean analyze the transferability of hand-crafted image features. They use a “nearest class multiple centroids” (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm. In combination with metric learning, they compare favorably against the $m$-nearest neighbor algorithm. Our work, on the other hand, highlights efficient transfer learning provided by foundation models.
@@
Using embedding vectors from pre-trained models is a simple and widely used transfer learning approach, established in the seminal works on computer vision (donahue_decaf_2014) and natural language processing (devlin_bert_2019). kornblith_better_2019 show that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks. In addition, huh_what_2016 provide insights into the required quality of pre-training data. Our work is orthogonal to these studies, since we focus on studying weight generation, normalization, and aggregation techniques applied later on for new task adaptation.
@@
Although we investigate transfer learning scenarios, we review the imprinting applications and results from CL. icarl dynamically select a subset of examples for each class and update internal representations via gradient descent. They use a nearest mean classifier (NMC) with respect to the saved examples. ncmbaseline use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings. However, they do not investigate the effect of normalization and using multiple proxies.
Findings of imprintallknn show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored. In our work, however, we compare imprinting all data to a limited number of more representative proxies striving for efficiency.
@@
The phenomenon of NC was identified by papyan2020prevalence and refers to the convergence of the last-layer weight vectors to class means. It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) (zhu2021geometric; neuralcollapse2; neuralcollapse3). Nevertheless, complete neural collapse is practically unrealistic (DBLP:conf/icml/TirerHN23). In transfer learning, DBLP:conf/iclr/Galanti0H22 show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios. In our work, we expand the survey on NC by experimenting with out-of-distribution classes belonging to different datasets and linking their degree of collapse to the success of certain imprinting strategies.
3 Methods
3.1 IMPRINT
In order to find out how to best and efficiently set the classifier weights of a foundational model FM in new, previously unseen tasks T, we create the IMPRINT framework (see fig.˜1) that generalizes previous work, specifically all the methods that work without access to cross-class statistics and gradient-based training. Thereby, we unify all the existing imprinting strategies described in section˜2.
@@ 3.1 IMPRINT
We analyze the effect of weight generation (GEN), normalizations ($\texttt{NORM}=\{\texttt{NORM}\textsubscript{{pre}},\texttt{NORM}\textsubscript{{post}},\texttt{NORM}\textsubscript{{inf}}\}$), and aggregation (AGG). The IMPRINT framework depicted in fig.˜1 consists of three main building blocks: a foundation model FM, a weight generator GEN, and extendable classifier weights that are imprinted. The FM remains frozen throughout the experiments. It receives data from T as inputs and produces embedding vectors. The training process generates weight vectors for each of the $C$ classes in T. Hereby, embeddings from the FM are normalized before the generation (GEN) step according to NORMpre. The generated weight vectors per class are referred to as proxies, prototypes, or representatives (metric-proxies; prototypical; convolutional-prototypes; semanticdriftCL). These proxies are normalized according to NORMpost. As in the work by qi-imprinting, we do not use bias values. To classify the test data in T during inference, it is first embedded by the FM, normalized according to NORMinf, and finally aggregated by AGG, resulting in a predicted class label.
Generalization of Previous Methods.
Previously proposed imprinting methods can be defined as special cases of our framework. We design IMPRINT such that every method can be defined by a single combination of GEN, NORM, and AGG, of which we inspect all possible combinations.
@@
The purpose of GEN is to determine how the embeddings of the training data in T are used to form the new weights. In contrast to qi-imprinting which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in first-nearestclassmean and enable non-linear classification. We denote the number of proxies per class as $k$, ranging between $1$ and the number of samples, and investigate the following operations conducted per class to generate those:
@@
  • k-means: $k$-means cluster centers using KMeans from sklearn (scikit-learn). $k=1$ is the same as mean.
  • k-medoids: $k$-medoids cluster centers using KMedoids from sklearn.
  • k-cov-max (covariance-maximization): Top $k$ embeddings by covariance (in code: proxies = embeddings[torch.argsort(torch.sum(torch.cov(embeddings), dim=0), descending=True)[:k]]).
  • k-fps (farthest-point sampling): Iteratively selecting $k$ embeddings, such that it maximizes the distance from already selected ones (starting with a random sample).
We choose this diverse list of methods to cover a wide range of approaches, ranging from heuristics (e.g., k-fps) to more complex algorithms (e.g., k-means). Note that only mean and k-means generate proxies beyond existing samples by producing synthetic cluster centers in embedding space. In comparison, k-medoids must choose actual samples as cluster centers (analogous to a median). None of these methods use cross-class statistics or gradient-based optimization. We also note an analogy to associative memory models, interpreting imprinting as a memory update process following a covariance rule, as detailed in section˜A.3.
@@
L2 normalization can be applied to embeddings before GEN via NORMpre, to the generated weights via NORMpost, and to embeddings in inference via NORMinf. In any case, the vector is $L^{2}$-normalized by dividing it by its Euclidean length $\|\cdot\|_{2}$.
quantile normalization (quantnorm1; quantnorm2) can only be applied to generated weights. This non-linear operation distributes weights equally. Recall that if more than one class is contained in T ($c>1$), GEN is performed for each class consecutively, and the reference distribution changes accordingly. In particular, for the first class there is no reference distribution to map to. This is different from itsdone, where new weights are matched to the distribution of the original classifier weights of the FM. Since we do not consider the classes used for pre-training the FM and especially do not assume access to their last-layer weights, this is not possible in our scenario.
@@
There are various ways to use the generated weights (proxies) per class during inference, especially when $k>1$. We focus on two different modes, max and $m$-nearest neighbor (m-nn). The former, max, computes the inner product of the input embedding and the imprinted weights and outputs the class label with the maximum activation. The latter, m-nn, uses the class weights as keys and the embeddings as values, and chooses the final winning output class via the $m$-nearest neighbor algorithm. The m-nn voting is weighted by the inverse of the Euclidean distances to their nearest neighbor, turning it into weighted majority voting.
We also experimented with alternative distance functions (Cosine, Manhattan) and uniform weighting, but found differences well within statistical noise. Using the Mahalanobis distance proved computationally prohibitive. More elaborate voting mechanisms, such as learned top-$k$ attention or entropy-based filtering, are conceivable, but extend well beyond the minimalist imprinting paradigm and therefore left for future work.
Note that max is the same as 1-nn in the case of L2 for NORMpost, since for any fixed embedding vector $\boldsymbol{\mathbf{v}}$ and variable proxy $\boldsymbol{\mathbf{w}}$, the argmin of $\|\boldsymbol{\mathbf{v}}-\boldsymbol{\mathbf{w}}\|^{2}=\|\boldsymbol{\mathbf{v}}\|^{2}-2\langle\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{w}}\rangle+\|\boldsymbol{\mathbf{w}}\|^{2}$, calculated by 1-nn, is the same as the argmax of the inner product $\langle\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{w}}\rangle$ calculated in max.
3.2 Quantifying Neural Collapse
Neural collapse (NC) (papyan2020prevalence) characterizes the state of the features produced by a classification neural network after training to near zero training loss. Namely, the learned embeddings of each class converge, i.e., collapse, to their class means. These globally centered class means and classifier weights form a simplex equiangular tight frame (ETF) – a collection of equal length and maximally equiangular vectors, that maximize the between-class variability. This results in an optimal linearly separable state for classification. In fig.˜3 (left), we illustrate the collapse of a FM on its pre-training data. The newly arrived data T from a different dataset is distributed more unevenly across the embedding space (right).